OpenStudy (anonymous):

Determine the number and type of solutions for x2 + 5x + 6 = 0. ANSWERS PLEASE

myininaya (myininaya):

\[b^2-4ac>0 => \text{ two real solutions}\] \[b^2-4ac<0 => \text{ two imaginary solutions}\] \[b^2-4ac=0 => \text{ one real solution with multplicity 2}\]

OpenStudy (anonymous):

in the form of \[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

RATIONAL OR IRATIONAL

OpenStudy (anonymous):

\[b^2-4ac \] \[5^2-4*1*6=1\] \[-b \pm \sqrt{\Lambda}/2a\] \[-5\pm1/2\] -4/2=-2 or -6/2=-3 one solution is -2 other is -3